OpenStudy (anonymous):

can somebody help me with this? Evaluate each polynomial for the given value of the variable. –x2-5x+6; x= -3

OpenStudy (welshfella):

replace all the x's in the polynomial with -3 and work it out;- -(-3)^2 - 5(-3) + 6 = ? notice its best to put the -3 in parentheses

OpenStudy (welshfella):

-5x means '-5 times x' so -5(-3) means -5 times -3

OpenStudy (anonymous):

how would I check my answer? is there a way or no

OpenStudy (anonymous):

the answer is 30 @welshfella ?

OpenStudy (anonymous):

9+15+6=30

OpenStudy (welshfella):

-(-3)^2 - 5(-3) + 6 = -9 + 15 + 6 = -9 +12 = ?

OpenStudy (anonymous):

isn't a negative times a negative a positive?

OpenStudy (welshfella):

-2 * -3 = 9 but there's a - sign before the x^2

OpenStudy (welshfella):

if the first term was (-x)^2 then it would be 9 but as its written its -x^2

OpenStudy (anonymous):

im so confused sorry

OpenStudy (anonymous):

-9 + 15 + 6 = -9 +12 = ? where did you get the 12?

OpenStudy (welshfella):

oh sorry its 15 + 6 = 21

OpenStudy (welshfella):

so that's -9 + 21

OpenStudy (welshfella):

a typo

OpenStudy (anonymous):

so the answer is 12?

OpenStudy (anonymous):

@welshfella